OpenStudy (dmezzullo):

Check My Work? Type the first equation from the group you selected and identify the slope and the y-intercept of the equation. Describe how you would graph this line using the slope-intercept method. Be sure to write in complete sentences. Type the second equation from the group you selected and change the equation into slope-intercept form. Be sure to show all of your work. Identify the slope and y-intercept of the equation. Describe how you would graph this line using the slope-intercept method. Be sure to write in complete sentences,

OpenStudy (dmezzullo):

In a paragraph of at least three sentences, explain how the graphs of the two equations in your group are the same and how they are different. Be sure to use complete sentences.

OpenStudy (dmezzullo):

2. y =-3 over 2 x + 4: slope =-3 y-intercept=4 3.to graph the line, put a point on the (0,4) and (-3,3) then just connect the lines and extend them. 4. –3x + 2y = 8 slope-intercept form: y=8-2x over -3 slope=2 y-intercept=8 5.put a point on (0,8) and (-2,7) then connect lines and extend. 6. They have different slopes . They have different y intercepts to. They Are Both Linear.

OpenStudy (dmezzullo):

^^^

OpenStudy (dmezzullo):

@LoveYou*69 @Hero

OpenStudy (dmezzullo):

wat u mean?

OpenStudy (dmezzullo):

ohhh ok.

OpenStudy (anonymous):

i thought u put at least 5 sentneces or u get pended

OpenStudy (anonymous):

@dmezzullo

OpenStudy (dmezzullo):

huh?

OpenStudy (anonymous):

nothin

OpenStudy (dmezzullo):

ok?